Farttex 1-6 marks) For the date given below; 6, 0, 2,-1, 5,6 I, The expected value of x is 2. The variance of s is 3. Ihe median is (0) 3 th) () 6 (d) 5 (a) 167 (b) 7.8 (e) 2.79 (d) 3.2 (b) 4 (c) 3.5 (d) 8 4. The mode is (a) 4 The range isg (a) S (h) H (c) 3 (d) 6 (h) H (c) 7 (d) 6 Panta· (2 mark) For o normal population with known variance ?. , the confidence level for the interval x 1.645a/n is; 7 (a) 0.95 (b) 0.025 (c) 0.05 (d) 0.90
Answer #1

Answer to the question is:

A. The expected value is basically the average of the values here : 6+ 0+ 2-1+5+6/6 = 3

B. Variance is basically the following :

((6-3)^2 + (0-3)^2+(2-3)^2+(-1-3)^2+(5-3)^2+(6-3)^2)/6 = 8, the nearest answer is 7.8

Please note that there is answer matching 8.

c. The median is the Middlemore value. When there are even nmbeen of observations we take out the mean of middle values. These are 2 and 5, which 3.5.

Answer : median is 3.5

4. Mode is basically the highest occuring number. 6 sports twice, so it's the mode.

Answer is 6

5. Range basically is the difference beteeen min and max

So range is 6-(-1)= 7

Range is 7